OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

me neither, although both are easy enough to solve

OpenStudy (anonymous):

lol ok ill try factoring

OpenStudy (anonymous):

\(y = x^2 + 2x + 1=(x+1)^2\) so as a perfect square it has the vertex on the \(x\) axis at \((-1,0)\)

OpenStudy (anonymous):

\(y = x^2 + 3x + 2=(x+1)(x+2)\) and so it crosses the \(x\) axis at \((-1,0)\) and \((-2,0)\) not sure what you are supposed to say about what they have in common, they are both parabolas that open up

OpenStudy (anonymous):

ok thank youu so much

OpenStudy (anonymous):

yw